OpenStudy (anonymous):

Can somebody help me memorize squares of 1-20? I need to memorize by tomorrow.

OpenStudy (anonymous):

There is no way I could help you to much. I would say make some flash cards. Do this. Put the number on one side and the square on the other and try to do this. Look at the number and guess the square then look on the back and see if you got it right. Retry on the ones you didn't

OpenStudy (anonymous):

Hope this helps.

zepdrix (zepdrix):

Oh boy :O The squares are important numbers. I can't think of any cool shortcut methods though hmm

OpenStudy (anonymous):

Thank you for trying to help me. I will certainly try this method, but I may be unable to do so as this is a school issue iPad and it might not let me.

OpenStudy (anonymous):

At least you try.

OpenStudy (anonymous):

At least. Thanks again!

OpenStudy (anonymous):

Yup np.

OpenStudy (anonymous):

Sorry, I do not know what np means please explain

OpenStudy (zzr0ck3r):

np = no problem

OpenStudy (anonymous):

^

OpenStudy (anonymous):

Oh. Thanks! I am a noob at this and at social media :\

OpenStudy (anonymous):

Its more internet slang. But not many people understand

OpenStudy (zzr0ck3r):

5*5=25 is just one bigger than 6*4 6*6=36 is just one bigger than 7*5 7*7=49 is just one bigger than 8*6 . . .

OpenStudy (zzr0ck3r):

this may come in handy...

OpenStudy (anonymous):

Oh... That's really cool! Thanks for helping, but I need to do mental math, including the ones above ten... Thanks again though!

OpenStudy (beginnersmind):

I'm guessing you know your multiplication tables up to 10x10. So the problem is 11 through 19. It really depends on how good your mental arithmetic is, sometimes you can just do it in your head. My favourite method is to think of something like 17*17 as (10 + 7) * (10 + 7) = 100 + 2*70 + 49 = 289. All the terms are fairly easy, especially if you have the square of 1 through 10 down cold. It also serves as a mnemonic (an aide to remember or recall some fact). I think it might help if you wrote out every square from 11 squared to 19 squared like this. After all, learning square numbers shouldn't just be about memorizing a fact. These numbers come from somewhere and going through the work of calculating them might help you remember as well.